OpenStudy (anonymous):

3x^2+18=0 10x^2-11x-6=0

hero (hero):

3x^2 + 18 = 0 3(x^2 + 6) = 0 x^2 + 6 = 0 x^2 = -6 x = ±sqrt(-6) x = ±sqrt(6)i

hero (hero):

For the second problem, use quadratic formula a = 10, b = -11, c = -6 \[x = \frac{-b ±\sqrt{b^2 - 4ac}}{2a}\]